### v1

# Distinguishing Power of 4-Legendrian Permutation Racks

Abstract
We explore 4-Legendrian rack structures and the effectiveness of 4-Legendrian racks to distinguish Legendrian knots. We prove that permutation racks with a 4-Legendrain rack structure cannot distinguish sets of Legendrian knots with the same knot type, Thurston–Bennequin number, and rotation number.
KEYWORDS: 4-Legendrian rack, invariant, Legendrian knot, Legendrian rack, permutation rack, rotation number, Thurston–Bennequin number
1 Introduction
Joyce and Matveev both independently introduced quandles to the mathematical world in the 1980s [8, 13]. A few years later, Fenn and Rourke popularized a generalization of quandles, which they called “racks” [5]. These racks were first used to study framed knots and links, but it was quickly realized that every semi-framed non-split link embedded in a 3-manifold has a fundamental rack that classifies both the link and the 3-manifold.
In 2017, Kulkarni and Prathamesh [12] had the idea to use rack invariants to study Legendrian knots. In 2021, Ceniceros, Elhamdadi, and Nelson [1] generalized this into a more powerful structure called a Legendrian rack. In 2023, Kimura [10] and Karmakar, Saraf, and Singh [9] further independantly generalized these structures into the more powerful GL-rack, also called a generalized Legendrian rack or a bi-Legendrian rack. In 2024, Kimura [11] even further generalized GL-racks into 4-Legendrian racks, a potentially more powerful structure and the object of study for this paper.
In Section 2, we review common rack theory ideas and definitions and provide several examples of families of racks that we explore later in this paper.
In Section 3, we review the history of Legendrian racks up to this point, including many recent developments.
In Section 4, we define 4-Legendrian racks and discuss their classification as algebraic structures.
In Section 5, we review classical and rack-theoretic invariants of Legendrian knots.
In Section 6, we prove our main theorem, which is as follows.
Theorem 1.1 (Theorem 6.3).
4-Legendrian permutation racks cannot distinguish isotopy classes of Legendrian knots with the same classical invariants.
Kimura [10] and Cheng and He [2] both posed the open question of whether there exist 4-Legendrian racks that distinguish Legendrian knots sharing the same topological knot type and classical invariants. This question remains open in general. However, since permutation racks are the main examples of racks that are not quandles in the literature, our main theorem and a similar result of Kimura for 4-Legendrian quandles [*][Thm. 4.2.3]kimura-thesis settle this question in the negative for extremely large classes of 4-Legendrian racks.
2 Review of Racks
Definition 2.1 (Rack, [3]).
A rack, $(X,\triangleright)$, is a set X with binary operation $\triangleright:X\times X\to X$ satisfying:
  1. 1.
    R1 (Invertibility): For all y $\in$ X, the map $\beta_{y}:X\to X$ defined by $\beta_{y}(x)=x\triangleright y$ is invertible. We denote $\beta^{-1}_{y}(x)$ as $x\triangleright^{-1}y$.
  2. 2.
    R2 (Self-Distributivity): For all x, y, z $\in$ X, $(x\triangleright y)\triangleright z=(x\triangleright z)\triangleright(y\triangleright z)$.
R1 and R2 are known as the rack axioms. A rack is called a quandle when the binary operation is also idempotent, that is for all $x\in X,x\triangleright x=x.$ We say that $|X|$ is the order of $(X,\triangleright)$. Finite racks have been enumerated up to isomorphism for orders up to 13 and fully classified for orders up to 11 [18].
In the following, let $n\geq 0$ be a nonnegative integer.
Example 2.2 (Trivial quandle).
The trivial quandle of order $n$, denoted $T_{n}$, has underlying set $X=\mathbb{Z}/n\mathbb{Z}$ and quandle operation $x\triangleright y\coloneq x$.
Example 2.3 (Dihedral quandle).
The dihedral quandle of order $n$, denoted $R_{n}$, has underlying set $X=\mathbb{Z}/n\mathbb{Z}$ and quandle operation $x\triangleright y\coloneq 2y-x\pmod{n}$.
Example 2.4 (Alexander quandle).
Pick $t\in\mathbb{Z}/n\mathbb{Z}$ such that $\gcd(t,n)=1$. The quandle with underlying set $X=\mathbb{Z}/n\mathbb{Z}$ and quandle operation $x\triangleright y\coloneq(1-t)y+tx$ is called an Alexander quandle of order $n$.
Example 2.5 (Conjugation quandle).
Let $G$ be any group. The conjugation quandle is the quandle $\operatorname{Conj}(G)\coloneq(G,\triangleright)$ with operation $a\triangleright b\coloneq bab^{-1}$.
Example 2.6 (Core quandle).
Let $G$ be any group. The core quandle is the quandle $\operatorname{Core}(G)\coloneq(G,\triangleright)$ with operation $a\triangleright b\coloneq ba^{-1}b$.
Example 2.7 (Takasaki quandle).
Let $A$ be any additive abelian group. The Takasaki quandle is the quandle $T(A)\coloneq(A,\triangleright)$ with operation $a\triangleright b\coloneq 2b-a$.
Remark 2.8.
When $A$ is abelian, Core($A$) = $T(A)$. Additionally, choosing $A=\mathbb{Z}/n\mathbb{Z}$ for Takasaki quandles yields the dihedral quandle, $R_{n}$.
All of the above examples are quandles. Examples of racks that are not quandles are much less common in the literature; the following is the usual example.
Example 2.9 (Permutation rack).
Let $X$ be a set, and let $\sigma\in S_{X}$ be a permutation of $X$. The permutation rack or constant action rack is the rack $X_{\sigma}\coloneq(X,\triangleright)$ with operation $x\triangleright y\coloneq\sigma(x)$.
Remark 2.10.
Letting $\sigma\coloneq\operatorname{id}_{X}$ results in the trivial quandle. Otherwise, $X_{\sigma}$ is a rack that is not a quandle as it will violate idempotency for at least one element.
Example 2.11 ($(t,s)$-rack).
Let $t$ be an invertible variable, $s$ a non-invertible variable such that $s^{2}=s(1-t)$, and $X$ be a $\mathbb{Z}[t^{\pm},s]/(s^{2}-s(1-t))$-module. The rack with underlying set $X$ and rack operation $x\triangleright y\coloneq tx+sy$ is called a $(t,s)$-rack.
Remark 2.12.
All (t, 1-t)-racks are Alexander quandles. When $s\neq 1-t$, the resulting rack is not a quandle as $x\triangleright x\neq x$.
2.1 Rack automorphisms
Definition 2.13.
Let $(X,\triangleright_{X})$ and $(Y,\triangleright_{Y})$ be racks. A map $\varphi\colon X\to Y$ is called a rack homomorphism if $\varphi(x_{1}\triangleright_{X}x_{2})=\varphi(x_{1})\triangleright_{Y}\varphi(x_{2})$ for all $x_{1},x_{2}\in X$.
If $(X,\triangleright_{X})=(Y,\triangleright_{Y})$ and $\varphi$ is bijective, we say that $\varphi$ is a rack isomorphism. A rack isomorphism from $(X,\triangleright_{X})$ to itself is a rack automorphism. The automorphism group of $(X,\triangleright_{X})$ is denoted by $\operatorname{Aut}(X)$.
Important to the theory of racks is the following canonical automorphism $\pi$. The name comes from the map’s association with kinks in diagrams of framed knots.
Definition 2.14 ([3, p. 149]).
Let $(X,\triangleright)$ be a rack. The kink map $\pi\colon X\to X$ is the function defined by $x\mapsto x\triangleright x$.
It can be shown that $\pi$ is a rack automorphism that commutes with all other rack homomorphisms; that is, $\pi$ lies in the center of the category of racks. Moreover, the inverse map is given by $\pi^{-1}(x)=x\triangleright^{-1}x$. See, for example, [15, Sec. 2.2].
Example 2.15.
A rack $(X,\triangleright)$ is a quandle if and only if $\pi=\operatorname{id}_{X}$. Thus, $\pi$ can be thought of as the obstruction to being a quandle.
Example 2.16.
If $X_{\sigma}$ is a permutation rack, then $\pi=\sigma$.
Another important class of rack automorphisms is the following.
Definition 2.17.
Let $(X,\triangleright)$ be a rack. The inner automorphism group $\operatorname{Inn}(X)$ (also called the operator group or right multiplication group) is the subgroup of the symmetric group $S_{X}$ generated by the maps $\beta_{x}$ ranging over all $x\in X$:
$\operatorname{Inn}(X)\coloneq\langle\beta_{x}\mid x\in X\rangle.$
Remark 2.18.
The rack axioms state precisely that $\operatorname{Inn}(X)$ is a subgroup of $\operatorname{Aut}(X)$. A quick check shows that this subgroup is normal.
3 History of Legendrian Racks
Although rack theory [5] and Legendrian knot theory [4] were popularized concurrently in the 1990s, it was not until 2017 when Kulkarni and Prathamesh [12] introduced the first rack-theoretic invariants of Legendrian knots. They called these invariants $n$-Legendrian racks, and they used these invariants to distinguish infinitely many Legendrian unknots [12, Main Thm. 2]. In 2021, Ceniceros, Elhamdadi, and Nelson [1] generalized $n$-Legendrian racks by introducing Legendrian racks, which are 4-Legendrian racks where $u_{l}=u_{r}=d_{l}=d_{r}$. They used Legendrian racks to distinguish certain Legendrian trefoils and connected sums of Legendrian trefoils [1, Sec. 5].
Further generalizing Legendrian racks, Kimura [10] and Karmakar, Saraf, and Singh [9] independently introduced GL-racks (also called generalized Legendrian racks or bi-Legendrian racks) in 2023. GL-racks are 4-Legendrian racks where $u_{l}=u_{r}$. They distinguish infinitely many Legendrian unknots and Legendrian trefoils; see [10, Thm. 4.1] and [9, Thms. 4.7 and 4.8]. In 2025, Cheng and He [2, Thm. 1.1] further showed that GL-racks distinguish Legendrian knots at least up to the absolute values of their classical invariants.
However, none of the above examples of Legendrian knots answer the question of whether rack invariants can distinguish Legendrian knots with the same topological knot type and classical invariants. In fact, in his 2024 PhD thesis introducing 4-Legendrian racks, Kimura [11, Sec. 4.2] found several examples of nonequivalent Legendrian knots sharing these invariants that cannot be distinguished using 4-Legendrian racks. He also showed that 4-Legendrian quandles cannot distinguish any such Legendrian knots [11, Thm. 4.2.3]. Our main theorem (Theorem 6.3) extends this result to all 4-Legendrian permutation racks. It is worth noting that while Kimura’s result uses a theorem from contact topology, our approach is purely algebro-combinatorial.
On the algebraic side of things, the first author [15] in 2025 presented a simplified but equivalent characterization of GL-racks. This led to a group-theoretic classification of GL-structures, answering a question posed by Karmakar, Saraf, and Singh in a previous version of [9]. Combining [15, Thm. 5.6] and [16, Thm. 10.1] yields the surprising result that the categories of racks, Legendrian racks, and GL-quandles are isomorphic. Moreover, certain involutory GL-racks have connections to symmetric racks (also called racks with good involutions) [*][Thm. 10.1]Ta_2025goodinvolutionsconjugationsubquandles, which are used to distinguish surface-knots in $\mathbb{R}^{4}$.
4 4-Legendrian Racks
In the following, let $(X,\triangleright)$ be a rack.
Definition 4.1 (Cf. [15, Sec. 4.1]).
Define the group $U_{X}$ to be the centralizer
$U_{X}:=C_{\operatorname{Aut}(X)}(\operatorname{Inn}(X)).$
We say that elements of $U_{X}$ are GL-structures on $(X,\triangleright)$. Ordered pairs of GL-structures $(u_{l},u_{r})\in U_{X}\times U_{X}$ are called 4-Legendrian structures on $(X,\triangleright)$.
Note that, by Remark 2.18, $U_{X}$ is a normal subgroup of $\operatorname{Aut}(X)$.
Definition 4.2.
A 4-Legendrian rack is a quadruple $(X,\triangleright,u_{l},u_{r})$ where $(X,\triangleright)$ is a rack and $(u_{l},u_{r})$ is a 4-Legendrian structure on $(X,\triangleright)$.
Remark 4.3.
A GL-rack is precisely a 4-Legendrian rack in which $u_{l}=u_{r}$. A Legendrian rack is precisely a GL-rack in which $d_{l}=u_{r}$ and (as a result) $d_{r}=u_{l}$.
Remark 4.4.
Since $U_{X}$ is nonempty, every rack $(X,\triangleright)$ can be equipped with at least one 4-Legendrian structure. A similar statement holds for GL-racks but not for Legendrian racks.
Definition 4.5.
Let $(X,\triangleright_{X},u_{l},u_{r})$ and $(Y,\triangleright,v_{l},v_{r})$ be 4-Legendrian racks. A rack homomorphism (resp. isomorphism) $\varphi\colon X\to Y$ is called a 4-Legendrian rack homomorphism (resp. isomorphism) if $\varphi u_{l}=v_{l}\varphi$ and $\varphi u_{r}=v_{r}\varphi$.
Example 4.6 (Cf. [10, Ex. 3.7]).
Let $X_{\sigma}$ be a permutation rack. Then a 4-Legendrian structure is precisely a pair $(u_{l},u_{r})$ of permutations of $X$ that commute with $\sigma$.
Example 4.7 (Cf. [10, Ex. 3.6]).
Let $G$ be a group, and let $g,h\in Z(G)$. Let $u_{l},u_{r}\colon G\to G$ be the multiplication maps $x\mapsto gx$ and $x\mapsto hx$. Then $(\operatorname{Conj}(G),u_{l},u_{r})$ is a 4-Legendrian quandle.
Remark 4.8.
In his PhD thesis introducing 4-Legendrian racks, Kimura [11, Sec. 4.2] defined a 4-Legendrian rack as a sextuple $(X,\triangleright,u_{l},u_{r},d_{l},d_{r})$ in which $(X,\triangleright)$ is a rack and $u_{l},u_{r},d_{l},d_{r}\colon X\to X$ are functions that satisfy the following eight axioms for all $x,y\in X$:
$d_{l}u_{r}=u_{r}d_{l}$ $=d_{r}u_{l}=u_{l}d_{r},$ (4.1)
$d_{r}u_{l}(x\triangleright x)$ $=x,$
$d_{l}(x\triangleright y)$ $=d_{l}(x)\triangleright y,$
$u_{l}(x\triangleright y)$ $=u_{l}(x)\triangleright y,$
$d_{r}(x\triangleright y)$ $=d_{r}(x)\triangleright y,$
$u_{r}(x\triangleright y)$ $=u_{r}(x)\triangleright y,$
$x\triangleright d_{l}(y)$ $=x\triangleright y=x\triangleright d_{r}(y),$
$x\triangleright d_{l}(y)$ $=x\triangleright y=x\triangleright d_{r}(y).$
Morphisms in this category are defined in the obvious way.
With some work, one can show that Kimura’s definition of 4-Legendrian racks is equivalent to our definition; $d_{l}$ and $d_{r}$ are respectively determined entirely by $u_{r}$ and $u_{l}$ as well as $\triangleright$:
$d_{l}=u_{r}^{-1}\pi^{-1},\qquad d_{r}=u_{l}^{-1}\pi^{-1}.$ (4.2)
In fact, the forgetful functor defined by $(X,\triangleright,u_{l},u_{r},d_{l},d_{r})\mapsto(X,\triangleright,u_{l},u_{r})$ is an isomorphism of categories. The proof of this fact is similar to the analogous result for GL-racks [15, Prop. 3.12]; we omit the details here.
In light of this equivalence, we use our definition while studying algebraic properties of 4-Legendrian racks, and we use Kimura’s definition while studying Legendrian knot invariants.
Example 4.9.
Given a 4-Legendrian permutation rack as in Example 4.6, we have
$(u_{r}d_{l})^{-1}=\sigma=(u_{l}d_{r})^{-1}.$
4.1 Classification of 4-Legendrian Racks
Let $(u_{l},u_{r})$ and $(v_{l},v_{r})$ be 4-Legendrian structures on $(X,\triangleright)$. By definition, the 4-Legendrian racks $(X,\triangleright,u_{l},u_{r})$ and $(X,\triangleright,v_{l},v_{r})$ are isomorphic if and only if there exists a rack automorphism $\varphi\in\operatorname{Aut}(X)$ such that $v_{l}=\varphi u_{l}\varphi^{-1}$ and $v_{r}=\varphi u_{r}\varphi^{-1}$. Equivalently, the pairs $(u_{l},u_{r})$ and $(v_{l},v_{r})$ are simultaneously conjugate in $\operatorname{Aut}R$; that is, the 4-Legendrian structures lie in the same orbit of $U_{X}\times U_{X}$ under the diagonal conjugation action of $\operatorname{Aut}(X)$. This action exists because $U_{X}$ is a normal subgroup of $\operatorname{Aut}(X)$. To summarize, we have just shown the following.
Proposition 4.10 (Cf. [15, Thm. 4.1]).
The isomorphism classes of 4-Legendrian structures on $(X,\triangleright)$ are precisely the orbits of $U_{X}\times U_{X}$ under the diagonal conjugation action of $\operatorname{Aut}(X)$.
Example 4.11.
Let $n\geq 0$ be a nonnegative integer, and let $T_{n}$ be the trivial quandle of order $n$. Then $U_{X}=\operatorname{Aut}(T_{n})=S_{n}$, so isomorphism classes of $4$-Legendrian racks with underlying rack $T_{n}$ correspond to orbits of $S_{n}\times S_{n}$ under the diagonal action of $S_{n}$ by conjugation. These orbits are counted in OEIS sequence A110143 [7]; see [17]. (We verified this example for all $n\leq 6$ using the computer search described below.)
Remark 4.12.
Even if $(X,\triangleright,u_{l})\cong(X,\triangleright,v_{l})$ and $(X,\triangleright,u_{r})\cong(X,\triangleright,v_{r})$ as GL-racks, the 4-Legendrian racks $(X,\triangleright,u_{l},u_{r})$ and $(X,\triangleright,v_{l},v_{r})$ are not necessarily isomorphic. For example, let $T_{3}$ be the trivial quandle of order $3$, and let $u_{l}=u_{r}=v_{l}=(2,3)$ and $v_{r}=(1,3)$ in cycle notation.
Remark 4.13.
Given a 4-Legendrian rack $(X,\triangleright,u_{l},u_{r})$, it is usually not isomorphic to $(X,\triangleright,u_{r},u_{l})$. For example, let $u_{l}$ be any nonidentity GL-structure on a rack $(X,\triangleright)$, and let $u_{r}\coloneq\operatorname{id}_{X}$.
Using Proposition 4.10, we implemented a GAP [6] program that classifies all $4$-Legendrian racks of a given order $n\leq 11$ up to isomorphism. This program uses the classification of racks up to order $11$ from [18], which is where the $n\leq 11$ bound comes from.
We were able to complete the search for all $n\leq 6$; our code and data is available in a GitHub repository [14]. Table 1 enumerates our data. The table also includes an enumeration of involutory 4-Legendrian racks; this is motivated by the connection between involutory GL-racks and symmetric racks shown in [16, Sec. 10].
Order 0 1 2 3 4 5 6
Racks 1 1 8 33 249 1592 15944
Involutory racks 1 1 8 24 196 850 9248
Quandles 1 1 4 16 84 448 3137
Kei 1 1 4 16 74 342 2228
Table 1: Number of $4$-Legendrian structures on various families of racks of order up to 66, counted up to isomorphism.
Corollary 4.13.1.
Let $G\coloneq\operatorname{Aut}(X)$. If $\operatorname{Inn}(X)\leq Z(G)$, then the isomorphism classes of 4-Legendrian structures on $(X,\triangleright)$ are precisely the orbits of $G\times G$ under the diagonal conjugation action of $G$. In particular, if $G$ is abelian, then these isomorphism classes are precisely the group $G\times G$.
Proof.
This follows from the fact that $\operatorname{Inn}(X)\leq Z(G)$ if and only if $U_{X}=G$. ∎
Example 4.14.
Let $n\in\mathbb{Z}^{+}$ be a positive integer, let $\sigma\in S_{n}$ be an $n$-cycle, and let $X_{\sigma}$ be the corresponding permutation rack of order $n$. Then $\operatorname{Aut}(X_{\sigma})=\langle\sigma\rangle\cong\mathbb{Z}/n\mathbb{Z}$, which is abelian, so the set of isomorphism classes of 4-Legendrian structures on $R$ is $\langle\sigma\rangle\times\langle\sigma\rangle$. In particular, there are exactly $n^{2}$ isomorphism classes of 4-Legendrian racks with underlying rack $X_{\sigma}$. (Our GAP search verified this fact for all $n\leq 6$.)
Example 4.15.
Let $F$ be the free rack on one generator, identified as the permutation rack $\mathbb{Z}_{\sigma}$ defined by $\sigma(k):=k+1$ for all $k\in\mathbb{Z}$. Then $\operatorname{Aut}(F)=\langle\sigma\rangle\cong\mathbb{Z}$, which is abelian. Hence, the set of isomorphism classes of 4-Legendrian structures on $F$ is $\{(\sigma^{m},\sigma^{n})\mid m,n\in\mathbb{Z}\}\cong\mathbb{Z}^{2}$.
5 4-Legendrian Rack Invariants
5.1 Legendrian Knots
Legendrian knots are important objects of study in the field of contact topology. We briefly review several concepts from the theory; we refer the reader to the survey of Etnyre [4] for a rigorous treatment.
The standard contact structure is the kernel of the differential 1-form dzydxdz-y\,dx in $\mathbb{R}^{3}$, which is depicted in Figure 1. A smooth knot in $\mathbb{R}^{3}$ is called Legendrian if it lies everywhere tangent to the standard contact structure.
Figure 1: The standard contact structure on $\mathbb{R}^{3}$, depicted as an assignment of a plane to each point.
Legendrian knots are usually studied via their front projections to the $xz$-plane. Front projections of Legendrian knots have two key features that distinguish them from projections of smooth knots. First, since tangent lines can never be vertical, front projections have cusps in place of vertical tangencies. Second, due to the direction of twisting, the overstrand at each crossing is always the strand having the more negative slope. Note that an oriented front projection never has two leftward-oriented or rightward-oriented cusps placed adjacent to each other. For example, Figure 3 depicts a Legendrian left handed trefoil in its front projection.
A major problem in contact topology is the classification of Legendrian knots up to Legendrian isotopy. Every smooth knot type has infinitely many Legendrian representatives. Many of them can be distinguished using their topological knot type and their two classical invariants, which we define below. Given a front projection of a Legendrian knot $L$, let $w(L)$ denote the writhe of the front projection, and let $U$ and $D$ denote the numbers of upward-oriented and downward-oriented cusps. Note that $U$ and $D$ must be positive, and $U+D$ must be even.
Definition 5.1.
The Thurston–Bennequin number of $L$ is the integer
$\operatorname{tb}(L)\coloneq w(L)-\frac{1}{2}(D+U).$
The rotation number of $L$ is the integer
$\operatorname{rot}(L)\coloneq\frac{1}{2}(D-U).$
These two integers are called the classical invariants of $L$.
Example 5.2.
The classical invariants of the Legendrian left-handed trefoil in Figure 3 are $(\operatorname{tb},\operatorname{rot})=(-6,-1)$.
However, there exist many nonequivalent Legendrian knots that share the same topological knot type and classical invariants. This motivates the search for Legendrian knot invariants that distinguish such pairs of Legendrian knots.
Definition 5.3.
A Legendrian knot invariant is called effective if it distinguishes some pair of Legendrian knots sharing the same topological knot type and classical invariants.
In particular, Kimura [10] and Cheng and He [2] posed the open questions of whether GL-racks and 4-Legendrian racks provide effective invariants of Legendrian knots. While this question remains open in general, Kimura [11, Thm. 4.2.3] gave a negative answer for all 4-Legendrian quandles, and our main theorem (Theorem 6.3) provides a negative answer for all 4-Legendrian permutation racks.
5.2 Coloring Invariants
In [11, Sec. 4.2], Kimura introduced an invariant of Legendrian knots $L$ called the fundamental 4-Legendrian rack, which we denote by $F(L)$. This is defined in the same vein as the fundamental GL-rack, which Karmakar, Saraf, and Singh introduced in [9, Thm. 4.3], and the fundamental quandle of a smooth knot, which was independently introduced by Joyce [8] and Matveev [13].
We outline the construction here; see [11, Sec. 4.2] for details, and cf. [9, 2]. 4-Legendrian racks can be viewed as an algebraic theory with two binary operations $\triangleright^{\pm 1}$ and four unary operations $u_{l}^{\pm 1},u_{r}^{\pm 1}$. It follows from general results from universal algebra that we can consider free 4-Legendrian racks, which are defined using the usual universal property, and take quotients of 4-Legendrian racks by congruence relations.
Fix an oriented front projection of a Legendrian knot $L$. Label the arcs (i.e., connected components) of the front projection by $x_{1},\dots,x_{n}$, and let $F\coloneq\langle x_{1},\dots,x_{n}\rangle$ be the free 4-Legendrian rack generated by $x_{1},\dots,x_{n}$. Starting at any crossing in the front projection, traverse the knot using its given orientation. At each cusp, impose a label $u_{l},u_{r},d_{l}$, or $d_{r}$ in the obvious way, depending on the orientation of the cusp; cf. (4.2). At each crossing, impose a relation on $F$ between arcs as illustrated in Figure 2. Define $F(L)$ to be $F$ modulo the congruence relation generated by these $n$ crossing relations.
Two local pictures of crossings in front projections of Legendrian knots. The first is a negative crossing, and the second is a positive crossing. The arcs are labeled using the relations of the fundamental 4-Legendrian rack.
Figure 2: Relations imposed on $F(L)$ between arcs at negative and positive crossings.
It can be shown that the isomorphism class of $F(L)$ is an invariant of $L$; see [11, Sec. 4.2]. Of course, this invariant is difficult to use directly, so we instead consider colorings of $F(L)$ by finite 4-Legendrian racks. These are easier to compute since $F(L)$ is finitely presented (having $n$ generators and $n$ relations in the above construction).
Definition 5.4.
Let $R\coloneq(X,\triangleright,u_{l},u_{r})$ be a 4-Legendrian rack. A coloring of $L$ by $R$ is a 4-Legendrian rack homomorphism $F(L)\to R$.
Since $F(L)$ is an invariant of $L$, the set $\operatorname{Hom}(F(L),X)$ of colorings of $L$ by a given 4-Legendrian rack $(X,\triangleright,u_{l},u_{r})$ is also an invariant of $L$; see [11, Rem. 30].
Example 5.5.
Consider the Legendrian left handed trefoil with classical invariants $(\operatorname{tb},\operatorname{rot})=(-6,-1)$ in Figure 3. The fundamental 4-Legendrian rack is the quotient of the free 4-Legendrian rack $\langle x_{1},x_{2},x_{3}\rangle$ by the congruence relation generated by the relations
$x_{3}\triangleright x_{1}=d_{l}u_{r}(x_{2}),\quad x_{2}\triangleright x_{3}=u_{l}d_{r}(x_{1}),\quad x_{1}\triangleright x_{2}=u_{l}u_{r}(x_{3}).$
A front projection of a Legendrian left handed trefoil. The arcs are labeled in a way that corresponds to the fundamental 4-Legendrian rack.
Figure 3: Legendrian left handed trefoil with arcs labeled.
6 Main Theorem
Fix an oriented front projection of a Legendrian knot $L$ with $U$ upward-oriented cusps and $D$ downward-oriented cusps, and let $(X_{\sigma},u_{l},u_{r},d_{l},d_{r})$ be a permutation rack equipped with a 4-Legendrian structure. To prove our main result, we construct a canonical form for relations in the image of a given coloring of $L$ by $X_{\sigma}$.
This idea is similar to an approach Cheng and He took to prove the main theorem of [2]. Namely, they observed that every relation in the canonical presentation of the fundamental GL-rack of $L$ can be written in the form
$x$ $=u^{p_{n}}d^{q_{n}}(x_{n})\triangleright^{\epsilon_{n}}x_{k_{n}}$
$=u^{p_{n}+p_{n-1}}d^{q_{n}+q_{n-1}}(x_{n-1}\triangleright^{\epsilon_{n-1}}x_{k_{n}-1})\triangleright^{\epsilon_{n}}x_{k_{n}}$
$=\dots$
$=u^{U}d^{D}(\dots(x\triangleright^{\epsilon_{1}}x_{k_{1}})\triangleright^{\epsilon_{2}}\dots)\triangleright^{\epsilon_{n}}x_{k_{n}},$
where $U=\sum p_{n}$, $D=\sum q_{n}$, and $\epsilon_{n}\in\{\pm 1\}$ with $\sum\epsilon_{n}=w(L)$.
The obstruction to taking this exact approach with 4-Legendrian racks is that $u_{l}$ and $d_{r}$ generally do not commute with $u_{r}$ and $d_{l}$. Nevertheless, we have the following.
Proposition 6.1.
Let $x$ be one of the generators of $F(L)$, and consider either of the two relations of $F(L)$ corresponding to a crossing in which $x$ is an understrand. Then this relation can be written in the form
$x=W(\dots(x\triangleright^{\epsilon_{1}}x_{k_{1}})\triangleright^{\epsilon_{2}}\dots)\triangleright^{\epsilon_{n}}x_{k_{n}},$ (6.1)
where $W$ is a word of length $U+D$ consisting of $U$ letters in $\{u_{l},u_{r}\}$ and $D$ letters in $\{d_{l},d_{r}\}$ (without inverses), and $\epsilon_{n}\in\{\pm 1\}$ with $\sum\epsilon_{n}=w(L)$. Moreover, $W$ alternates between letters in $\{u_{l},d_{l}\}$ and letters in $\{u_{r},d_{r}\}$.
Proof.
In the given relation for $x$, it follows from (4.2) that we can move all of the cusp functions $u_{l},u_{r},d_{l},d_{r}$ to the leftmost part of the word, giving a string $W$ of length $U+D$.
To obtain the defining relations for $F(L)$, we had to traverse the front projection of $L$ using its given orientation. It follows that none of the relations obtained in this way contain the inverse of any of the cusp functions. Moreover, it is impossible for two rightward-oriented cusps or two leftward-oriented cusps to appear in a row. It follows that our word alternates between left and right cusp functions. Hence, $W$ satisfies the desired properties. The rest is identical to the above calculation of Cheng and He. ∎
Lemma 6.2.
In the image of any coloring of $L$ by a 4-Legendrian permutation rack $X_{\sigma}$, any word of the form (6.1) can be rewritten in either the form
$x=\underbrace{(d_{l}d_{r}\dots d_{l}d_{r})}_{\operatorname{rot}(L)\text{ times}}\sigma^{\operatorname{rot}(L)+\operatorname{tb}(L)}(x)\hskip 18.06749pt\text{or}\hskip 18.06749ptx=\underbrace{(d_{r}d_{l}\dots d_{r}d_{l})}_{\operatorname{rot}(L)\text{ times}}\sigma^{\operatorname{rot}(L)+\operatorname{tb}(L)}(x).$
Proof.
Since $X_{\sigma}$ is a permutation rack, the string to the right of $W$ in (6.1) collapses to $\sigma^{w(L)}$; that is, $x=W\sigma^{w(L)}$.
Since $L$ is a Legendrian knot, $U+D$ must be even, and $U,D>0$. It follows from Proposition 6.1 that one of $u_{l}d_{r}$, $d_{r}u_{l}$, $u_{r}d_{l}$, or $d_{l}u_{r}$ appears in $W$. Introduce a $\sigma\sigma^{-1}$ in the middle of this pair. Since $\sigma$ must commute with $u_{l},u_{r},d_{l},\text{ and }d_{r}$, we can commute the $\sigma^{-1}$ to the rightmost part of the word. By Example 4.9, we can then use the $\sigma$ to cancel out the $u_{l}d_{r}$, $d_{r}u_{l}$, $u_{r}d_{l}$, or $d_{l}u_{r}$. We can repeat this action until there are no longer any upward cusp functions in our word.
The word now only contains $D-U=2\operatorname{rot}(L)$ total $d_{l}$’s and $d_{r}$’s. Either $d_{l}$ or $d_{r}$ begin the word and the two functions must alternate. Meaning we are left with $\operatorname{rot}(L)$ pairs of $d_{l}d_{r}$’s or $d_{r}d_{l}$’s followed by $w(L)-U=\operatorname{rot}(L)+\operatorname{tb}(L)$ total $\sigma$’s in our word, as desired. ∎
We now prove the main theorem.
Theorem 6.3 (Theorem 1.1).
44-Legendrian permutation racks cannot distinguish isotopy classes of Legendrian knots with the same classical invariants.
Proof.
Let $(X_{\sigma},u_{l},u_{r},d_{l},d_{r})$ be an any 44-Legendrian permutation rack. Assume $L_{1}$ and $L_{2}$ are two distinct Legendrian isotopy classes of knots with the same classical invariants. Fix oriented front projections of $L_{1}$ and $L_{2}$.
In the image of a coloring of $L_{1}$ or $L_{2}$ by $(X_{\sigma},u_{l},u_{r},d_{l},d_{r})$, we have $x\triangleright y=\sigma(x)$ at each crossing relation. Therefore, we can combine all the relations for $L_{1}$ and $L_{2}$ into a single generator and single relation $\langle x:x=R_{L_{1}}(x)\rangle$ and $\langle x:x=R_{L_{2}}(x)\rangle$, where $x$ can start at any arc of the front projections of $L_{1}$ and $L_{2}$, respectively.
By Lemma 6.2, both of the relations $R_{L_{i}}(x)$ can be written in either the form
$x=\underbrace{(d_{l}d_{r}\dots d_{l}d_{r})}_{\operatorname{rot}(L)\text{ times}}\sigma^{\operatorname{rot}(L)+\operatorname{tb}(L)}(x)\hskip 18.06749pt\text{or}\hskip 18.06749ptx=\underbrace{(d_{r}d_{l}\dots d_{r}d_{l})}_{\operatorname{rot}(L)\text{ times}}\sigma^{\operatorname{rot}(L)+\operatorname{tb}(L)}(x).$
We can choose x in each respective front projection such that the word made of $d_{l}$’s and $d_{r}$’s begins with the same function. Since $L_{1}$ and $L_{2}$ share the same classical invariants, it follows that these two relations are identical. Hence, the number of homomorphisms from the fundamental 44-Legendrian rack to $(X_{\sigma},u_{l},u_{r},d_{l},d_{r})$ will be equal for both $L_{1}$ and $L_{2}$. ∎
Acknowledgments
We would like to thank the organizers of the Unknot V conference for giving the authors a chance to meet each other and Jose Ceniceros, whose work with Mohamed Elhamdadi and Sam Nelson we are building off of. The second author would also like to thank Lenny Ng, Roger Casals, Orsola Capovilla-Searle and Tye Lidman for fruitful conversations concerning Legendrian knots.